OpenStudy (firejay5):

Solve each system of equations by using substitution: The Algebra 2 Math Book says #17 is a no solution. How is it show your work and steps on how it's not a solution. 17. 6x - 3y = 12 2x = 8 - y

OpenStudy (mathlegend):

So first you would have to get one of the variables by itself. Since the second equation has a single 'y' why not use that one?

OpenStudy (mathlegend):

We can add y to both sides then subtract 2x from both sides leaving us with... y=-2x+8

OpenStudy (mathlegend):

So now we can substitute y into the first equation.

OpenStudy (mathlegend):

6x-3(-2x+8)=12

OpenStudy (mathlegend):

6x+6x-24=12

OpenStudy (mathlegend):

12x-24=12

OpenStudy (mathlegend):

12x=36

OpenStudy (mathlegend):

x=3

OpenStudy (firejay5):

wait

OpenStudy (mathlegend):

What don't you understand?

OpenStudy (firejay5):

I am trying to get caught up

OpenStudy (mathlegend):

oh ok

OpenStudy (firejay5):

???

OpenStudy (mathlegend):

2x = 8 - y +y +y ---------- y+2x=8 -2x -2x --------- y=-2x+8 (Got it?)

OpenStudy (firejay5):

ready for next step

OpenStudy (mathlegend):

Okay

OpenStudy (mathlegend):

So now that we solved for x... we can substitute it into the 1st equation.

OpenStudy (mathlegend):

6(3) - 3y = 12

OpenStudy (mathlegend):

18 - 3y=12 -18 -18 ----------- -3y= -6 -------- y=2

OpenStudy (firejay5):

the book says it's not a solution

OpenStudy (mathlegend):

Yes it is a solution. So I do not know why your book says no. My guess would be a typo.

OpenStudy (firejay5):

okay

OpenStudy (mathlegend):

I also just plugged it into an equation checker and it shows it as a solution

OpenStudy (firejay5):

okay help me on my other problem, the people there are poop